# WanJuan-CC: A Safe and High-Quality Open-sourced English Webtext Dataset

### 05 Mar 2024 v4

@@ 3.1 Processing Pipeline
We arrange the order of the pipeline based on computational cost and iteration frequency. Steps with lower computational costs are prioritized, while those with higher iteration frequencies are placed later. For instance, heuristic rule filtering has a relatively low computational cost and can effectively filter out a large amount of invalid text, so we place it as the first step after text extraction. Despite the high computational load of global deduplication, its low iteration frequency leads us to position it as the second step. Both safety filtering and quality filtering have high computational costs, but the threshold for quality filtering may vary depending on the use case, while the threshold for safety filtering is relatively fixed. Therefore, we decide to perform safety filtering first, followed by quality filtering.
Next we will introduce each step’s specific implementation.
@@ 3.3 Heuristic Rule Filtering
Given the prevalence of parsing errors, formatting issues, and non-natural text in the extracted data, the typical solution is to design heuristic rules to modify and filter these data. Building on the heuristic rules established by Gopher[19] and C4[20], we have added some new rules based on our in-depth observation of the data. Firstly, we implement those rules that would alter the content of the data:
  1. 1.
[…]
  2. 2.
    Effective Line Extraction: Find the first and last lines containing punctuation, retain the content between these two lines, and discard the rest. If the first line contains more than three words, it is retained.
  3. 3.
    Line Content Modification: Remove certain patterns, such as: “cookie policy”, “use of cookies”, etc.
  4. 4.
    Line Content Removal: Delete lines with the following characteristics: only contain spaces; contain “lorem ipsum”, “javascript”, “{”, etc.
After applying the above rules to truncate and modify the documents, we execute the following rules for filtering:
  1. 1.
    Drop the doc with the most common character not being a letter;
  2. 2.
    Drop the doc with the ratio of letters to numbers falling below 0.46;
  3. 3.
    Drop the doc with the proportion of the most frequent word in a single document exceeding 7.5% of the total word count (for texts with a total word count of 500 or less, this threshold is 30%);
  4. 4.
    Drop the doc with the word count outside the range of 50 to 100000;
  5. 5.
    Drop the doc with the number of words containing at least one letter accounting for less than 80% of the total word count;
  6. 6.
    Drop the doc with fewer than two stop words;
  7. 7.
    Drop the doc with the average word length outside the range of 3 to 10 characters;
  8. 8.
    Drop the doc with fewer than three lines, or where the third longest line contains fewer than 20 characters;
  9. 9.
    Apply the n-gram duplicate filter proposed by Gopher [19];
  10. 10.
    Drop the doc with the ratio of the total character count to the total word count exceeding 1.2;
  11. 11.
[…]
  12. 12.
    Drop the doc with either a sequence of 500 consecutive spaces or 8 consecutive newline characters;
  13. 13.
    Drop the doc with the proportion of newline characters exceeding 25% of the total character count;
  14. 14.
    Drop the doc with special characters resulting from decoding errors;
  15. 15.
    Drop the doc with any word exceeding 45 letters in length;
  16. 16.
    Drop the doc with any sentence exceeding 56 words in length;
  17. 17.
    Drop the doc if the last character is a colon;
  18. 18.
    Drop the doc containing image URL or no content other than the URL.
After the document has been intercepted, modified, and filtered according to the above rules, we obtained the Clean Data.
@@ 3.4 Deduplication with MinHash-LSH
Specifically, we utilized the datasketch library[22] for computation, applying 128 hash functions on the 5-gram of the documents to construct signatures, and setting 0.7 as the deduplication threshold. During the processing, we prioritized retaining the most recent data, meaning that Common Crawl dumps with larger numbers were considered first. After the deduplication process using LSH (Locality-Sensitive Hashing), we obtained the Dedup data.
@@ 3.5.1 Block Domain/Word List
We created a blocklist of approximately 13 million unsafe domain names, along with a blocklist of 36,289 unsafe words. We employed domain matching and text matching methods to label all the data. Considering that unsafe words could potentially lead to excessive data deletion, we strived to minimize words that might cause such deletions. By using domain and word filters, we labeled a total of 7.4% of the data.
@@ 3.5.2 Toxicity Classification
To address the potential harmful content on the Internet, we utilized the training and test sets provided by the Toxic Comment Classification Challenge on the Kaggle platform[16] to fine-tune the Bert model for toxic data filtering. This dataset includes labels for six categories: toxic, severe_toxic, obscene, threat, insult, and identity_hate. Each piece of data has six 0/1 attribute tags, where 1 indicates that the data possesses the corresponding category attribute, and 0 indicates the absence thereof. We re-labeled each piece of data according to the following rule: if any of the six attributes is 1, it is marked as "toxic"; otherwise, it is marked as "non-toxic". In this way, we obtained a binary classification dataset, with a ratio of approximately 1:2.5 between toxic and non-toxic data. During this process, we did not alter the original division of the training and test sets in the dataset.
We fine-tuned the bert for sequence classification model on the re-annotated training set, and ultimately, with a threshold set at 0.5, the model’s precision reached 81.56%, recall was 87.95%, and the F1 score was 84.63%. Utilizing this model, we conducted a comprehensive screening of the Dedup Data dataset, marking all data exceeding a threshold of 0.2 as to-be-discarded. Approximately 3.9% of the data was marked for discarding.
@@ 3.5.3 Pornography Classification
Firstly, we referred to the publicly available collection of pornographic website domains on GitHub at https://github.com/Bon-Appetit/porn-domains/tree/master, which contains a total of 228,254 domain names of pornographic websites. From the Dedup Data, we extracted 100,000 data entries based on this list as samples of pornographic website domains, and inversely extracted another 100,000 data entries as samples of non-pornographic website domains. These samples were drawn proportionally from 90 dumps according to the volume of data in each dump.
In order to obtain a higher quality dataset, we further utilized the pornography content detection interface of Perspective API to annotate these data. Data with scores exceeding 0.1 were classified as pornographic content, while those scoring below 0.1 were considered non-pornographic. Ultimately, we selected 30,000 pieces of pornographic/non-pornographic content data as the training set for fine-tuning the BERT model. From this, we extracted 4,000 pieces of pornographic/non-pornographic content data, which after manual annotation, served as the test set.
In our manually annotated test set, by setting the threshold to 0.3, the optimal model we trained exhibited a precision of 0.79, a recall of 0.991668, and an accuracy of 0.87. We used this model to filter the entire Dedup Data, marking all data exceeding a threshold of 0.2 as to be discarded. Approximately 7.3% of the data was marked in this way.
@@ 3.5.4 Personal Identifiable Information Masking
To safeguard user privacy, we employed regular expressions to mask potential Personally Identifiable Information (PII) present in the data. The types of PII that we masked include: IP addresses, passport numbers, GPS location data, personal email addresses, personal phone numbers, identification card numbers, bank account numbers, and physical addresses. For each type of PII, we used corresponding regular expressions for matching, recorded their positions in the text, and then replaced them with “[[pii_type]]”.
@@ 3.6 Data Quality Filtering
After completing the aforementioned filtering steps, we found that the data still contained a large amount of promotional information and content with poor language fluency. The promotional information mainly manifested as enticing titles and content, which contributed minimally to the training effectiveness of the model. Content with poor language fluency included format errors, machine-generated random text, aggregation of irrelevant content, and content with extremely low information content. To eliminate these elements, we employed a Bert-based classifier for promotional content and language fluency for further data filtering.
@@ 3.6.1 Advertisement Classification
In dealing with advertisement attributes, we established a detailed manual annotation guide, requiring annotators to categorize the data into: a) full-text advertisements; b) partial advertisements; c) non-advertisement full text; d) invalid data. Here, "invalid data" represents data that cannot be effectively judged, and this portion of data will be excluded. For the remaining data, we tagged "full-text advertisements" and "partial advertisements" with advertisement content labels, and labeled "non-advertisement full text" with non-advertisement content labels, thereby constructing a binary classification dataset.
@@ 3.6.2 Fluency Classification
For the attribute of fluency, we constructed a more comprehensive manual annotation questionnaire, requiring annotators to evaluate from the following four data usability dimensions:
  1. 1.
    Consistency, that is, checking whether the data revolves around the same theme throughout, to avoid topic confusion caused by extraction errors.
  2. 2.
    Noisy, that is, judging whether there are a large number of meaningless characters in the data, such as garbled text, special symbols, excessive spaces, etc.
  3. 3.
    Information, that is, assessing whether the data contains valuable information, for example, whether the information is incomplete due to the lack of image or video data.
  4. 4.
    Grammar, that is, checking whether there are grammatical errors in the data, such as spelling mistakes, structural errors, improper use of punctuation marks, etc.
We normalize the weighted sum of scores across all dimensions for each data point. Data with a score greater than 0.5 is labeled as having good fluency, while data with a score less than 0.5 is marked as having poor fluency, thereby forming a binary classification dataset.
@@ 3.7 Data Quality Evaluation
After processing with the aforementioned data processing pipeline, we obtain a High-Quality Data. The design of rules and selection of filtering thresholds are manually evaluated and adjusted to achieve the desired effect. To ensure data quality, we have implemented a comprehensive data quality evaluation and feedback mechanism in our engineering practice. It is widely recognized that assessing the quality of the Common Crawl dataset presents significant challenges, primarily due to its large volume, diverse problem types, vague quality definitions, high degree of subjectivity, and scattered features.
Informed by industry standards and best practices, we assess the dataset employing seven primary quality evaluation metrics, each encompassing multiple secondary metrics, as delineated below:
  • Quality Signal: Effectiveness
    • Ratio of data items of null data values in the dataset.
    • Ratio of data items containing garbled characters in the dataset.
  • Quality Signal: Completeness
    • Ratio of data items in the dataset that contain incomplete content (such as ending with a colon).
    • Ratio of data items in the dataset where paired content does not appear in pairs (such as parentheses).
  • Quality Signal: Understandability
    • Ratio of data items in the dataset that contain errors in formatting (such as formulas, tables, markdown, etc.).
    • Ratio of data items in the dataset that contain special characters (such as high proportion of spaces, line breaks, #, invisible characters).
  • Quality Signal: Similarity
    • Ratio of data items in the dataset that contain duplicate content.
    • Ratio of duplicate data items in the dataset.
  • Quality Signal: Fluency
    • Ratio of data items in the dataset that contain large blocks of text without punctuation.
    • Ratio of data items in the dataset that contain anti-scraping content.
  • Quality Signal: Relevance
    • Ratio of data items in the dataset that contain leading or trailing off-topic/unrelated content (such as advertisements, copyright notices, references from web pages).
  • Quality Signal: Security
    • Ratio of data items in the dataset that contain pornography content.
    • Ratio of data items in the dataset that contain toxic content.
In order to conduct a comprehensive, accurate, and efficient data quality evaluation, we conceptualize the process into three distinct stages: metric identification and definition, computation methods for these metrics, and determination of metric thresholds. To address these stages, we employ three disparate quality assessment methodologies: 1) Human Evaluation Method; 2) Quality Signal Evaluation Method; 3) Model Evaluation Method.
Human Evaluation Method: Initially, given the large volume of Common Crawl data, ambiguous quality definitions, high subjectivity, and dispersed features, we extract a certain amount of data for multiple evaluators to identify ’bad samples’. Quality indicators are then abstracted from the characteristics of these identified bad samples.
Subsequently, in order to enhance the likelihood of encountering bad samples during sampling, we randomly select a specific quantity (1 million) of data points to calculate the loss via LLMs. Based on the distribution of high, medium, and low loss values, we designate a certain number of data points (2000 each from high and low intervals, 1000 from the middle interval) for further human evaluation.
Quality Signal Evaluation Method: To accurately and efficiently gauge the extent of data quality issues within the entire Common Crawl dataset, we have devised machine rules that are independent of the data cleaning stage. These rules automatically evaluate the proportion of poor-quality samples in the entire dataset based on the quality indicators identified through the Human Evaluation Method. We define this proportion as $X=\frac{A}{B}$, where $A$ represents the number of poor-quality samples, and $B$ denotes the total number of samples in the dataset.
However, not all quality indicators can yield an accurate proportion of the entire dataset via machine rules, especially for complex issues such as texts containing leading or trailing promotional information. For these types of quality indicators, we estimate the proportion of poor-quality samples as $\tilde{X}=\frac{A^{\prime}}{B^{\prime}}$, where $A^{\prime}$ is the number of poor-quality samples identified by Human Evaluation, and $B^{\prime}$ is the total number of sampled instances.
Model Evaluation Method: We have obtained the proportions of quality indicators through both human and machine methods, which serve to measure the severity of issues. However, determining what proportion is deemed serious or not requires an independent threshold for each quality indicator. We initially set these thresholds based on our experience, then update the signals and thresholds in response to feedback from downstream Model Evaluation. Firstly, we will use the 1B and 3B experimental models trained in Section 4.4 for evaluation. In the Model Evaluation phase, we have designed a test dataset comprising approximately 600 prompts, spanning over 20 types in both Chinese and English. We then conduct a comprehensive assessment of the model-generated content using both subjective evaluation and Quality Signal Evaluation. This process allows us to explore the impact of data quality issues on model performance, thereby facilitating updates to the quality indicators and thresholds within our quality standards.

### 18 Mar 2024 v6

@@ 3.3 Heuristic Rule Filtering
Given the prevalence of parsing errors, formatting issues, and non-natural text in the extracted data, the typical solution is to design heuristic rules to modify and filter these data. Building on the heuristic rules established by Gopher[19] and C4[20], we have added some new rules based on our in-depth observation of the data. Firstly, we implement a set of rules that aim to normalize and simplify the content of the data, including character normalization, effective line extraction, line content modification, and line content removal. After applying these rules to truncate and modify the documents, we execute a series of heuristic rules to filter the entire document. For example, the ratio of letters in the text is too low, there are too many spaces and line breaks, there is a lack of punctuation, or too long words, etc. After the document has been intercepted, modified, and filtered according to the above rules, we obtained the Clean Data.
@@ 3.4 Deduplication with MinHash-LSH
Specifically, we utilized the datasketch library[22] for computation, applying 128 hash functions on the 5-gram of the documents to construct signatures, and setting a certain threshold to identify highly similar data as duplicates. During the processing, we prioritized retaining the most recent data, meaning that Common Crawl dumps with larger numbers were considered first. After the deduplication process using LSH (Locality-Sensitive Hashing), we obtained the Dedup data.
@@ 3.5.1 Block Domain/Word List
We created a blocklist of approximately 13 million unsafe domain names, along with a blocklist of 36,289 unsafe words. We employed domain matching and text matching methods to label all the data. Considering that unsafe words could potentially lead to excessive data deletion, we strived to minimize words that might cause such deletions.
@@ 3.5.2 Toxicity Classification
To address the potential harmful content on the Internet, we utilized the training and test sets provided by the Toxic Comment Classification Challenge on the Kaggle platform[16] to fine-tune the Bert model for toxic data filtering. This dataset includes labels for six categories: toxic, severe_toxic, obscene, threat, insult, and identity_hate. We simplified these labels into a binary classification of "toxic" and "non-toxic", based on whether any of the original six attributes were present. In this way, we obtained a binary classification dataset. During this process, we did not alter the original division of the training and test sets in the dataset and fine-tuned the Bert model.
@@ 3.5.3 Pornography Classification
We utilized a publicly accessible list of pornographic website domains to extract a sample of 100,000 entries from our data. An equal number of non-pornographic samples were also extracted. These samples were proportionally drawn from various data dumps. To enhance the accuracy of our annotations, we employed an API for further annotation. This process resulted in a refined dataset of 60,000 items. We then selected a subset of this data for manual annotation as the test set, and used the remaining data to train and fine-tune our model.
@@ 3.5.4 Personal Identifiable Information Masking
To safeguard user privacy, we employed regular expressions to mask potential Personally Identifiable Information (PII) present in the data. The types of PII that we masked include: IP addresses, passport numbers, GPS location data, personal email addresses, personal phone numbers, identification card numbers, bank account numbers, and physical addresses. We used specific patterns for each type of PII, identified their positions in the text, and replaced them with a generic placeholder.
@@ 3.6 Data Quality Filtering
After the initial filtering, we noticed that the data still included a significant amount of promotional content and poorly written text. To address this, we used a Bert-based classifier to further filter the data based on these two aspects.
@@ 3.6.1 Advertisement Classification
In dealing with advertisement attributes, we established a detailed manual annotation guide, requiring annotators to categorize the data into advertisements or not advertisements. thereby constructing a binary classification dataset, and divided into training and test sets.
@@ 3.6.2 Fluency Classification
For the attribute of fluency, we constructed a more comprehensive manual annotation questionnaire, requiring annotators to evaluate the data based on several dimensions including consistency, noise level, information value, and grammar correctness. We assign an overall tag(fluent or non-fluent) to each piece of data based on scores over all dimensions, thereby constructing a binary classification dataset, and divided into training and test sets.
@@ 3.7 Data Quality Evaluation
After applying our data processing pipeline, we obtain a High-Quality Data. To ensure the quality of this data, we have implemented a comprehensive evaluation and feedback mechanism, taking into account the challenges posed by the large volume and diverse nature of the Common Crawl dataset.
In line with industry standards and best practices, we assess the dataset using a variety of quality evaluation metrics. These metrics cover several aspects such as effectiveness, completeness, understandability, similarity, fluency, relevance, and security. Each aspect includes multiple secondary metrics to provide a comprehensive evaluation.
To conduct a comprehensive, accurate, and efficient data quality evaluation, we employ three distinct methodologies: Human Evaluation, Quality Signal Evaluation, and Model Evaluation. These methods help us identify and measure the severity of data quality issues, and refine our evaluation and processing methods based on feedback from downstream model evaluations.